# A Polylogarithmic Competitive Algorithm for Stochastic Online Sorting and TSP

### v2

@@ 2.2 Probability of failure
Note that sometimes our algorithm might fail. In this section we will upper bound the probability of failure of our algorithm. There are two cases where our algorithm could fail. The first one is at line 10 of Algorithm˜1 and the other is at line 23. To show that the failure probability $\mathbb{P}\!\left[\text{fail}\right]$ is small, we first present the following lemma, which is central to the analysis of the balls-into-bins instances arising in our problem.
@@
Suppose a balls-into-bins instance with $K$ total bins, each with capacity $C=\Theta(\log^{2}n)$ and let $M=K\cdot C\leq n$ the total number of balls that all bins can collectively hold. If the probability that an element belongs to each bin is the same, then the first overflow of a bin happens after $M-\frac{M}{\Theta(\log^{1/2}n)}$ balls are thrown w.p. at least $1-K/n^{2}$. Also, after throwing $M+\frac{M}{\Theta(\log^{1/2}n)}$ balls, all bins are full w.p. at least $1-K/n^{2}$.
@@
  • $\mathcal{G}_{i}$: each bin of the instance has received at least $C-\frac{C}{\Theta(\log^{1/3}n)}$ balls after $M_{i}-\frac{M_{i}}{\Theta(\log^{1/2}n)}$ balls have been thrown in this instance
@@
We now proceed by handling the second case of failure (Algorithm˜4). First we introduce some definitions.
@@
Note that for phase $i$, we can write its balls-into-bins instance size as $M_{i}=A_{i}+N_{i-1}$
@@
For each phase $i$, define as $T_{i}$ the number of elements inserted during phase $i$ until the first overflow in $A_{i}$ and $T_{i}^{\prime}$ as the number of elements inserted during phase $i$ until $A_{i}$ becomes full.
@@
$T_{2}-(T_{1}^{\prime}-T_{1})$ $\geq M_{2}-\frac{M_{2}}{\Theta(\log^{1/2}n)}-\left(M_{1}+\frac{M_{1}}{\Theta(\log^{1/2}n)}-A_{1}+N_{1}\right)$ (3)
$\geq A_{2}+N_{1}-\frac{A_{2}+N_{1}}{\Theta(\log^{1/2}n)}-\left(A_{1}+\frac{A_{1}}{\Theta(\log^{1/2}n)}-A_{1}+N_{1}\right)$ (4)
$=A_{2}-\frac{A_{2}+N_{1}}{\Theta(\log^{1/2}n)}-\frac{A_{1}}{\Theta(\log^{1/2}n)}$ (5)
$>A_{1}/4-o(A_{1})\geq 0,$ (6)
@@
Induction step: Assume the statement holds for each $i\in[r-1]$. We aim to prove that it also holds for $i=r$. From induction hypothesis, since we condition on $\mathcal{E}$, every subarray $A_{j}$ for $j\in[r-1]$ becomes completely filled before any bucket in $A_{j+1}$ overflows. As a result, when phase $r+1$ starts, every subarray $A_{j}$ for $j\in[r-1]$ is filled. This is crucial, since it implies that every new sample that arrives is inserted in this balls-into-bins instance thus we can use Definition˜2.4 (otherwise the elements inserted into the instance are not necessarily uniform in bins). The size of the balls-into-bins instance of phase $r+1$ is $A_{r+1}+N_{r}$. Now, using Definition˜2.4 again, observe that:
@@
We now prove that $\mathcal{E}$ is a very high probability event.
@@
[…]
The proof of the lemma is deferred to the appendix. Since $\mathcal{E}$ implies that our algorithm will not fail we get that our algorithms does not fail with high probability since:
@@ 3 Stochastic Online Sorting
  • InBucketPlacement(x,a)(x,a): We use the deterministic adversarial algorithm $\mathcal{A}_{adv}$ of [AABK23]
@@
We begin by conditioning on the success of our algorithm, which ensures that each bucket in every subarray stores exactly the elements belonging to its designated interval. Specifically, when the last subarray overflows (since it has one bucket this means it is full) all the other subarrays are completely filled, thus we are done. We distinguish three sources of cost: the intra-bucket cost, the inter-bucket cost, and the cost of connecting different subarrays.
@@
  • Inter-bucket: Consider subarray $A_{i}$. Since, by definition, each bucket contains elements strictly smaller than the next, the total cost to "connect them" is at most $\sum_{j=1}^{K_{i}-1}\frac{j+1}{K_{i}}-\frac{j-1}{K_{i}}\leq 2$, thus in total for all subarrays: $O(\log n)$.
  • Between subarrays: Between subarrays we incur cost at most $1$ and since we have $k=O(\log n)$ transitions this cost is $O(\log n)$.
@@
As a result, by conditioning on the success of our algorithm, we obtain total cost $O(\log^{2}n)$, thus with probability at least 11/n1-1/n, concluding the proof of the theorem. ∎
@@
Our algorithm is conceptually simple to state and analyse, and benefits from a hierarchical decomposition of Stochastic Online Sorting into balls-into-bins instances. The algorithm and the decomposition above can be naturally extended to non-uniform distributions. The only additional step that we need to take care of, is to partition the $[0,1]$ interval into as many intervals as required in each phase so that there is equal probability that a new point arrives in each interval. Furthermore, by a more careful analysis and parameter selection, we can show an upper bound of $O(\log^{3/2+\epsilon}n)$ on the total cost of the algorithm w.h.p. for Stochastic Online Sorting. Finally, by choosing a sufficiently large constant $c$ in the proof of Lemma˜2.3, we obtain an even stronger high-probability guarantee, e.g. $1-1/n^{100}$.
@@ 4.1 DomainPartitioning
Figure 2: A visual representation of our splitting procedure on the plane. Uniform distribution ensures that splitting a block in half creates two block of equal probability mass.
We partition the domain into hyper-boxes, which we call blocks. Let $\ell$ be the unique positive integer such that $\frac{n}{4\log^{2}n}<2^{\ell}\leq\frac{n}{2\log^{2}n}.$ As in the one-dimensional case, the final block covers the entire domain, i.e., $T^{(\ell)}_{1}=\mathcal{S}$. To construct the partition of $T^{(i)}$, we apply the split procedure from Algorithm˜2 to $T^{(i+1)}$. Each split doubles the number of blocks, so after $\ell$ splits, $T^{(1)}$ consists of $2^{\ell}$ blocks. The splits proceed cyclically over the dimensions: the first along the first dimension, the second along the second, and so forth. After $d$ splits, we return to the first dimension and continue this process until all $\ell$ splits have been performed.
@@ 3   Split $b$ into two equal halves $b_{1},b_{2}$ along dimension $i$;
4 $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\{b_{1},b_{2}\}$;
@@
We refer to the blocks $T^{(1)}=\{b_{1},b_{2},\ldots,b_{2^{\ell}}\}$ at the lowest level of our hierarchy as elementary blocks. Within an elementary block, all points are treated identically by Algorithm˜1: by construction, the algorithm’s decisions do not depend on which specific point from the block is presented. Consequently, we may assume that the points drawn within an elementary block $b$ are i.i.d. samples from the uniform distribution $\mathcal{U}(b)$ [^2] Notice, $\mathcal{U}(b)$ is the restriction of the global distribution $\mathcal{D}$ to block $b$. .
@@
2
3while $\neg(v[1]=1\wedge\dots\wedge v[d-1]=1\wedge v[d]=n_{d})$ do
@@ 6    $j\leftarrow j+1$;
7
 $v[j]\leftarrow v[j]+m[j]$ ;
 /* $v$ now has the coordinates of the next block */
@@ 4.2 InBucketPlacement
Once an element is mapped to a bucket, it must be assigned to one of the empty cells within that bucket. For this step we rely on an existing result: we apply Bertram’s adversarial algorithm [Ber25] to determine the placement.
@@ 4.3 Cost Analysis
Given sufficiently large $n$, the expected length of a TSP tour with $n$ points drawn i.i.d from $\mathcal{U}([0,\Delta]^{d})$ is $\beta_{d}\cdot n^{1-1/d}\cdot\Delta$, where $\beta_{d}\approx\sqrt{\frac{d}{2\pi e}}$. We denote this quantity by $\text{TSP}(n,d,\Delta)$.
@@
    For any $S\subseteq[0,\Delta]^{d}$, let $tour(S,n)$ denote the expected length of the optimal TSP tour on $n$ points drawn uniformly i.i.d. in $S$. Then, $tour(S,n)\;\leq\;\text{TSP}(n,d,\max_{x,y\in S}\|x-y\|_{\infty})\;\leq\;\text{TSP}(n,d,\Delta).$
@@
$\mathbb{P}\!\left[|T-\text{TSP}(n,d,\Delta)|\geq t\right]\leq 2\exp\left(-\frac{ct^{2}}{n^{1-2/d}\Delta^{2}}\right)$
@@
In contrast to the one-dimensional case, there is no closed-form expression for the optimal cost in higher dimensions. We therefore rely on asymptotic estimates and concentration bounds to control the value of Opt.
@@
A feasible tour inside $r(B^{(j)}_{k})$ is obtained by concatenating the optimal tours of $r(b_{k_{1}}),\dots,r(b_{k_{m}})$ and adding at most (r1)(r-1) connecting edges between adjacent elementary blocks. Summed over all phase $j$ blocks, the total number of such connectors is at most the number of elementary blocks. By Lemma˜4.15, each connector has length at most $2\beta_{d}\;n^{-1/d}\log^{2}n$, and by Corollary˜4.16 the total intra-subarray connection cost is deterministically bounded by $\beta_{d}\,n^{1-1/d}$. This yields the claimed inequality. ∎
@@
The cost of connecting all blocks within a subarray is deterministically bounded by $\beta_{d}\cdot n^{1-1/d}$.
@@
  • Intra-bucket: For each subarray $A_{j}$, placements inside buckets are handled by Bertram’s algorithm [Ber25]. This incurs cost $\sum_{i=1}^{K_{j}}\sqrt{C_{i}}\cdot\textsc{Opt}(r(B_{i}))\leq\Theta(\log n)\sum_{i=1}^{2^{\ell}}\textsc{Opt}_{i}=\Theta(\log n)\textsc{Opt}$. Where the first inequality is derived Lemma˜4.9 and the second by Lemma˜4.7.Since there are $R=O(\log n)$ subarrays, the total bucket cost is $O(\log^{2}n)\textsc{Opt}$.
  • Inter-bucket: By Lemmas˜4.15 and 4.16, the additional cost of connecting all buckets within a subarray $A_{i}$ is at most $O(\textsc{Opt})$. Summing over all $k=O(\log n)$ subarrays yields a total of $O(\log n)\textsc{Opt}$.
  • Between subarrays: Connecting consecutive subarrays requires at most one additional edge per transition. By Lemma˜4.17, this incurs cost at most Opt per transition. With $k=O(\log n)$ subarrays, the total cost is $O(\log n)\textsc{Opt}$.
@@ 5 Conclusion and Open Directions
In this work we study the stochastic variants of Online Sorting and Online TSP, obtaining polylogarithmic upper bounds in both settings. Analyzing the problem through the lens of the balls-into-bins literature reveals exponential improvements over previous approaches. A natural direction is to extend our bounds to general known distributions in higher dimensions. At the same time, our framework also assumes full knowledge of the distribution. Relaxing this assumption prompts several questions: What guarantees are possible when points are drawn i.i.d. from an unknown distribution, or when inputs arrive in uniformly random order with no distributional assumptions?
@@ Appendix A Full Pseudocode
$\mathcal{T}\leftarrow\texttt{DomainPartitioning}\left(\mathcal{D},\mathcal{S},K\right)$ ;
@@ 12    $\texttt{idx}\leftarrow\texttt{idx}+c^{(i)}_{j}$
14
15 while $\forall j,\ a^{(i)}_{j}$ is not full do

### v3

@@ 2.2 Probability of failure
Note that sometimes our algorithm might fail. In this section we upper bound the probability of failure of our algorithm. There are two cases where our algorithm could fail. The first one is at line 10 of Algorithm˜1 and the other is at line 23. To show that the failure probability $\mathbb{P}\!\left[\text{fail}\right]$ is small, we first present the following lemma, which is central to the analysis of the balls-into-bins instances arising in our problem.
@@
Suppose a balls-into-bins instance with $K$ total bins, each with capacity $C=\Theta(\log^{2}n)$, and let $M=K\cdot C\leq n$ be the total number of balls that all bins can collectively hold. If the probability that an element belongs to each bin is the same, then the first overflow of a bin happens after $M-\frac{M}{\Theta(\log^{1/2}n)}$ balls are thrown w.p. at least $1-K/n^{2}$. Moreover, after throwing $M+\frac{M}{\Theta(\log^{1/2}n)}$ balls, all bins are full w.p. at least $1-K/n^{2}$.
@@
  • $\mathcal{G}_{i}$: each bin of the instance has received at least $C_{i}-\frac{C_{i}}{\Theta(\log^{1/3}n)}$ balls after $M_{i}-\frac{M_{i}}{\Theta(\log^{1/2}n)}$ balls have been thrown in this instance
@@
We now proceed by handling the second case of failure. First we introduce some definitions.
@@
Note that for phase $i$, the size of its balls-into-bins instance is $M_{i}=A_{i}+N_{i-1}$.
@@
For each phase $i$, define as $T_{i}$ the number of elements inserted during phase $i$ until the first overflow in $A_{i}$ and $T_{i}^{\prime}$ as the number of elements inserted since the beginning of phase $i$ until $A_{i}$ becomes full.
@@
$T_{2}-(T_{1}^{\prime}-T_{1})$ $\geq M_{2}-\frac{M_{2}}{\Theta(\log^{1/2}n)}-\left(M_{1}+\frac{M_{1}}{\Theta(\log^{1/2}n)}-A_{1}+N_{1}\right)$
$\geq A_{2}+N_{1}-\frac{A_{2}+N_{1}}{\Theta(\log^{1/2}n)}-\left(A_{1}+\frac{A_{1}}{\Theta(\log^{1/2}n)}-A_{1}+N_{1}\right)$
$=A_{2}-\frac{A_{2}+N_{1}}{\Theta(\log^{1/2}n)}-\frac{A_{1}}{\Theta(\log^{1/2}n)}$
$>A_{1}/4-o(A_{1})\geq 0,$
@@
Induction step: Assume the statement holds for each $i\in[r-1]$. We aim to prove that it also holds for $i=r$. From induction hypothesis, since we condition on $\mathcal{E}$, every subarray $A_{j}$ for $j\in[r-1]$ becomes completely filled before any bucket in $A_{j+1}$ overflows. As a result, when phase $r+1$ starts, every subarray $A_{j}$ for $j\in[r-1]$ is filled. This is crucial, since it implies that every new sample that arrives is inserted in this balls-into-bins instance; thus we can use Definition˜2.4 (otherwise the elements inserted into the instance are not necessarily uniform in bins). The size of the balls-into-bins instance of phase $r+1$ is $A_{r+1}+N_{r}$. Now, using Definition˜2.4 again, observe that:
@@
We establish the following lemma, whose proof is deferred to the appendix.
@@
It holds that $\mathbb{P}\!\left[\neg\mathcal{E}\right]\leq\frac{1}{n}.$
Since $\mathcal{E}$ implies that our algorithm will not fail we get that our algorithm does not fail with high probability since:
@@ 3 Stochastic Online Sorting
  • InBucketPlacement(x,a)(x,a): We use the deterministic adversarial algorithm $\mathcal{A}_{adv}$ of [AAB+23] to place elements within each bucket.
@@
We begin by conditioning on the success of our algorithm, which ensures that each bucket in every subarray stores exactly the elements belonging to its designated interval. We distinguish four sources of cost: the intra-bucket cost, the inter-bucket cost, the cost of connecting different subarrays and the cost of the final subarray $B$.
@@
  • Inter-bucket: Consider subarray $A_{i}$. Since, by definition, each bucket contains elements strictly smaller than the next, the total cost to "connect them" is at most $\sum_{j=1}^{K_{i}-1}\left(\frac{j+1}{K_{i}}-\frac{j-1}{K_{i}}\right)\leq 2$, thus in total for all subarrays: $O(\log n)$.
  • Between subarrays: Between subarrays we incur cost at most $1$ and since we have $R=O(\log n)$ transitions this cost is $O(\log n)$.
@@
As a result, by conditioning on the success of our algorithm, we obtain total cost $O(\log^{2}n)$, thus the total cost is $O(\log^{2}n)$ with probability at least 11/n1-1/n, concluding the proof of the theorem.
@@
Our algorithm is conceptually simple to state and analyze, and benefits from a hierarchical decomposition of Stochastic Online Sorting into balls-into-bins instances. The algorithm and the decomposition above can be naturally extended to non-uniform distributions. The only additional step we need to take care of is to partition the $[0,1]$ interval into as many intervals as required in each phase so that there is equal probability that a new point arrives in each interval. Furthermore, by a more careful analysis and parameter selection, we can show an upper bound of $O(\log^{3/2+\epsilon}n)$ on the total cost of the algorithm w.h.p. for Stochastic Online Sorting. Finally, by choosing a sufficiently large constant $c$ in the proof of Lemma˜2.3, we obtain an even stronger high-probability guarantee, e.g. $1-1/n^{100}$.
@@ 4.1 DomainPartitioning
Figure 2: A visual representation of our splitting procedure on the plane. Uniform distribution ensures that splitting a block in half creates two blocks of equal probability mass.
We partition the domain into hyperboxes, which we call blocks. Let $\ell$ be the unique positive integer such that $\frac{n}{4\log^{2}n}<2^{\ell}\leq\frac{n}{2\log^{2}n}.$ As in the one-dimensional case, the final block covers the entire domain, i.e., $T^{(\ell)}_{1}=\mathcal{S}$. To construct the partition of $T^{(i)}$, we apply the split procedure from Algorithm˜2 to $T^{(i+1)}$. Each split doubles the number of blocks, so after $\ell$ splits, $T^{(1)}$ consists of $2^{\ell}$ blocks. The splits proceed cyclically over the dimensions: the first along the first dimension, the second along the second, and so forth. After $d$ splits, we return to the first dimension and continue this process until all $\ell$ splits have been performed.
@@ 3   Split $b$ into two equal halves $b_{1},b_{2}$ along dimension $i$;
4 $\mathcal{B}^{\prime}\leftarrow\mathcal{B}^{\prime}\cup\{b_{1},b_{2}\}$;
@@
We refer to the blocks $T^{(1)}=\{b_{1},b_{2},\ldots,b_{2^{\ell}}\}$ at the lowest level of our hierarchy as elementary blocks. Within an elementary block, all points are treated identically by Algorithm˜1: by construction, the algorithm’s decisions do not depend on which specific point from the block is presented. Consequently, we may assume that the points drawn within an elementary block $b$ are i.i.d. samples from the uniform distribution $\mathcal{U}(b)$ [^2] Note that $\mathcal{U}(b)$ is the restriction of the global distribution $\mathcal{D}$ to block $b$..
@@
1
2while $\neg(v[1]=1\wedge\dots\wedge v[d-1]=1\wedge v[d]=n_{d})$ do
@@ 5    $j\leftarrow j+1$;
6
7 $v[j]\leftarrow v[j]+m[j]$, Append $v$ to $L$;
@@ 4.2 InBucketPlacement
Once mapped to a bucket, an element is placed in an empty cell using Bertram’s adversarial algorithm [BER25].
@@ 4.3 Cost Analysis
Given sufficiently large $n$, the expected length of a TSP tour with $n$ points drawn i.i.d. from $\mathcal{U}([0,\Delta]^{d})$ is $\beta_{d}\cdot n^{1-1/d}\cdot\Delta$, where $\beta_{d}\approx\sqrt{\frac{d}{2\pi e}}$. We denote this quantity by $\text{TSP}(n,d,\Delta)$.
@@
    For any convex $S\subseteq[0,\Delta]^{d}$, let $tour(S,n)$ denote the expected length of the optimal TSP tour on $n$ points drawn uniformly i.i.d. in $S$. Then, $tour(S,n)\;\leq\;\text{TSP}(n,d,\max_{x,y\in S}\|x-y\|_{\infty})\;\leq\;\text{TSP}(n,d,\Delta).$
@@
$\mathbb{P}\!\left[\left|T-\text{TSP}(n,d,\Delta)\right|\geq t\right]\leq 2\exp\left(-\frac{ct^{2}}{n^{1-2/d}\Delta^{2}}\right)$
@@
In contrast to the one-dimensional case, there is no simple closed-form expression for the optimal cost in higher dimensions. We therefore rely on asymptotic estimates and concentration bounds to control the value of Opt.
@@
A feasible tour inside $r(B^{(j)}_{k})$ is obtained by concatenating the optimal tours of $r(b_{k_{1}}),\dots,r(b_{k_{m}})$ and adding at most (r1)(r-1) connecting edges between adjacent elementary blocks. Summed over all phase $j$ blocks, the total number of such connectors is at most the number of elementary blocks. By Lemma˜4.15, each connector has length at most $2\beta_{d}\;n^{-1/d}\log^{2}n$, and by Corollary˜4.17 the total intra-subarray connection cost is deterministically bounded by $\beta_{d}\,n^{1-1/d}$. This yields the claimed inequality. ∎
@@
Remark 4.16.
Lemma 4.15 applies to elementary blocks. When moving to higher phases, blocks are formed by merging pairs of blocks from the previous phase. As a result, the diameter of a block (and hence the cost of connecting two neighboring blocks) increases by at most a factor strictly smaller than $2$ from one phase to the next. However, at the same time, the number of blocks to be connected decreases by a factor of $2$ at each phase. Therefore, the total cost incurred by connecting all blocks within a subarray remains of the same order, which leads to Corollary 4.17.
@@
Using the observation above, the cost of connecting all blocks within a subarray is deterministically bounded by $\beta_{d}\cdot n^{1-1/d}$.
@@
  • Intra-bucket: For each subarray $A_{j}$, placements inside buckets are handled by Bertram’s algorithm [BER25]. This incurs cost $\sum_{i=1}^{K_{j}}\sqrt{C_{j}}\cdot\textsc{Opt}(r(B_{i}))\leq\Theta(\log n)\sum_{i=1}^{2^{\ell}}\textsc{Opt}_{i}=\Theta(\log n)\textsc{Opt}$. where the first inequality is derived from Lemma˜4.9 and the second from Lemma˜4.7. Since there are $R=O(\log n)$ subarrays, the total bucket cost is $O(\log^{2}n)\textsc{Opt}$.
  • Inter-bucket: By Lemmas˜4.15 and 4.17, the additional cost of connecting all buckets within a subarray $A_{i}$ is at most $O(\textsc{Opt})$. Summing over all $k=O(\log n)$ subarrays yields a total of $O(\log n)\textsc{Opt}$.
  • Between subarrays: Connecting consecutive subarrays requires at most one additional edge per transition. By Lemma˜4.18, this incurs cost at most Opt per transition. With $k=O(\log n)$ subarrays, the total cost is $O(\log n)\textsc{Opt}$.
@@ 5 Conclusion and Open Directions
In this work, we study the stochastic variants of Online Sorting and Online TSP, obtaining polylogarithmic upper bounds in both settings. Analyzing the problem through the lens of the balls-into-bins paradigm reveals exponential improvements over previous approaches. A natural direction is to extend our bounds to general known distributions in higher dimensions. At the same time, our framework also assumes full knowledge of the distribution. Relaxing this assumption prompts several questions: What guarantees are possible when points are drawn i.i.d. from an unknown distribution, or when inputs arrive in uniformly random order with no distributional assumptions?
@@ Appendix A Full Pseudocode
$\mathcal{T}\leftarrow\texttt{DomainPartitioning}\left(\mathcal{D},\mathcal{S},\ell\right)$ ;
@@ 14    $\texttt{idx}\leftarrow\texttt{idx}+c^{(i)}_{j}$
16
17 while $\forall j\in[K_{i}],\ a^{(i)}_{j}$ is not full do